# A Systematic Literature Review of Software Engineering Research on Jupyter Notebook

### v1

@@ 6 Future Research Directions
As discussed in Section 5.1.2, existing code search techniques in Jupyter Notebook primarily use keyword matching or deep learning techniques to search the code directly based on natural language queries [186, 84]. However, these methods can be further improved by incorporating natural language summaries of code functionality into the search process. Future research should explore leveraging LLMs to automate the summarization of individual notebook code blocks, producing concise and descriptive explanations of their purposes. Such summaries not only enrich search accuracy, but also aid users in quickly assessing retrieved results without having to read and interpret the code directly.
@@
In Section 5.3.1, we discuss different code refactoring tools that offer basic functionalities for notebooks, such as splitting cells [146] and reordering notebook cells [159]. Future research should expand the capabilities of these refactoring tools, addressing common patterns such as the distribution of exploratory workflows across multiple notebooks [135]. In addition, researchers can target other refactoring features such as detecting and eliminating duplicate code, extracting reusable functions, and restructuring notebooks by merging cells to improve the readability and maintainability of notebooks.
@@
Research Direction 3: Generating Documentation for Groups of Dependent Code Cells
@@
Existing methods for automated documentation generation in notebooks (discussed in Section 5.4) focus primarily on the generation of documentation at the individual cell level. For example, Cell2Doc generates Markdown descriptions for single cells [100], and HeaderGen generates structural headings into the notebook [167]. However, these single-cell approaches often fail to capture broader contexts or overarching objectives spanning multiple interrelated cells that perform tasks collectively. Since notebook users intentionally cluster related operations into adjacent groups of cells to boost productivity and organization [183], future research should aim to automatically generate context-aware documentation for these clusters. Utilizing analysis techniques such as dependency graph extraction and runtime tracing, researchers can identify groups of cells and automatically produce Markdown summaries.
@@
Research Direction 4: Designing Automated Bug Detection and Remediation Tools Tailored to Notebooks
@@
Since bug analysis is crucial to maintaining the quality of Jupyter notebooks (see Section 5.5), future research could focus on developing automated tools specifically designed for notebook-related bugs. Although some existing approaches address issues such as name-value inconsistencies detection [109] and data leakage detection [197], further research is needed to tackle more complex bugs, such as cell defects and kernel failure. Future studies should focus on addressing notebook code cell defects, which arise from missing dependencies, incorrect execution order, or undefined variables, as well as kernel crashes, often caused by excessive resource consumption, infinite loops, or faulty package interactions [26]. Advanced techniques, including static and dynamic code analysis, execution flow tracking, and visualization-based debugging, can be leveraged to efficiently detect, categorize, and resolve these bugs.
@@
Research Direction 5: Conducting Empirical Studies on the Impact of Best Practices for Coding
@@
As outlined in Section 5.7, little is known about the actual influence of following best practices for coding in Jupyter notebooks. Future studies should conduct empirical studies to quantify the effects of following best practices for coding based on categorized effective and ineffective practices [116]. Systematic studies comparing compliant and non-compliant notebooks could provide concrete empirical evidence highlighting the advantages of rigorous adherence to recommended practices. Such evidence would strongly motivate wider adoption of disciplined coding standards among notebook users.
@@
Research Direction 6: Detecting and Resolving Duplicate Execution Numbers in Notebook Cells
@@
Research Direction 7: Enhancing AI-Powered Code Generation and Workflow Assistance in Jupyter Notebook
In Section 5.9, we discuss existing solutions for generating code snippets based on natural language instructions for specific libraries like Pandas [201]. These solutions possess limited scope with a single library and fail to support broader, multi-library, multi-step data science workflow automation. To address these limitations, future research should leverage advanced AI techniques (e.g., large language models), to improve conversational AI-powered assistants for automated code generation by multi-step workflow automation. For example, interactive low-code panels that recommend follow-up analytical questions and visualize workflow structures [22] could be refined and augmented through continuous conversational AI interactions. By dynamically adapting their recommendations based on real-time user inputs and evolving contexts, these AI-powered assistants could offer more responsive and intuitive exploratory data analysis experiences for notebook users.
@@ 7 Threats to Validity
External validity: While most studies analyzed in this review focus on Python notebooks, which may limit generalizability, this focus is justified by the fact that 93% of notebooks available on GitHub are written in Python [63]. This high percentage highlights the relevance of Python-based research within the broader context of Jupyter Notebook usage. Another concern regarding external validity is that our findings primarily apply to software engineering research on Jupyter notebooks. Future research should expand the scope by investigating other types of research on notebooks to improve the generalizability of the results.
Construct validity: One potential threat to the construct validity of our review is related to the study selection process. We conducted a comprehensive search of all academic articles listed in the DBLP bibliography, intentionally excluding non-academic sources such as blog posts. The DBLP bibliography is widely recognized and frequently used by software engineering researchers to identify relevant articles in systematic literature reviews [193, 32, 82].
@@ 8 Conclusion
In this systematic literature review, we provide a comprehensive overview of the research landscape of software engineering in Jupyter notebooks. Our review uncovers critical insights and trends that have shaped the field’s current state. We utilized a thorough methodology involving identifying, screening, and analyzing 146 relevant research articles (supplementary data are available here [^14] https://zenodo.org/records/15226809). The most important findings of our study are:
@@
    Most studies were published in conference venues, indicating the field is rapidly evolving. Interestingly, the most frequent venues were HCI-related instead of core SE venues. The Conference on Human Factors in Computing Systems (CHI) published 15 studies, whereas the International Conference on Software Engineering (ICSE) published 6.
@@
    Notebook-specific solutions for software engineering issues, such as testing, refactoring, and documentation, are relatively underexplored. For example, resolving duplicate execution numbers, refactoring inter-notebook clones, and generating group documentation for coherent code cells are future directions derived from our study.
@@
    The majority of replication packages are hosted on GitHub, raising concerns about long-term availability and adherence to open science principles, as GitHub repositories can be volatile. In contrast, only 13 studies used permanent repositories such as Zenodo or Figshare, aligning better with open science best practices.

### v2

@@ 6 Future Research Directions
As discussed in Section 5.1.2, existing code search techniques in Jupyter Notebook primarily use keyword matching or deep learning techniques to search the code directly based on natural language queries [watson2019pysnippet, li2021nbsearch]. However, these methods can be further improved by incorporating natural language summaries of code functionality into the search process. Future research should explore leveraging LLMs to automate the summarization of individual notebook code blocks, producing concise and descriptive explanations of their purposes. Such summaries not only enrich search accuracy, but also aid users in quickly assessing retrieved results without having to read and interpret the code directly.
@@
In Section 5.3.1, we discuss different code refactoring tools that offer basic functionalities for notebooks, such as splitting cells [shankar2022bolt] and reordering notebook cells [titov2022resplit]. Future research should expand the capabilities of these refactoring tools, addressing common patterns such as the distribution of exploratory workflows across multiple notebooks [rule2018exploration]. In addition, researchers can target other refactoring features such as detecting and eliminating duplicate code, extracting reusable functions, and restructuring notebooks by merging cells to improve the readability and maintainability of notebooks.
Research Direction 3: Inclusive and Accessible Computational Notebooks
Despite the widespread adoption of computational notebooks, accessibility remains a largely underexplored research area (discussed in Section 5.3.3). Prior studies have shown that accessibility remains a critical challenge in computational notebooks, with large-scale analyses revealing widespread structural barriers that limit usability for blind and visually impaired users [venkatesh2023notably]. In particular, many notebooks lack essential semantic elements, such as headings and alternative text for visualizations, and rarely provide accessible representations of graphical outputs. While initial tools like MatplotAlt enable authors to add alternative text to figures [nylund2025matplotalt], future research should move beyond isolated solutions and develop accessibility-aware notebook environments that support inclusive documentation by default. This includes automated detection of accessibility violations, generation of alternative text and accessible tables alongside visual outputs, and integration of accessibility guidance directly into notebook authoring workflows to ensure that accessibility is treated as a first-class quality attribute of computational notebooks.
@@
Research Direction 4: Generating Documentation for Groups of Dependent Code Cells
@@
Existing methods for automated documentation generation in notebooks (discussed in Section 5.4) focus primarily on the generation of documentation at the individual cell level. For example, Cell2Doc generates Markdown descriptions for single cells [mondal2023cell2doc], and HeaderGen generates structural headings into the notebook [venkatesh2023enhancing]. However, these single-cell approaches often fail to capture broader contexts or overarching objectives spanning multiple interrelated cells that perform tasks collectively. Since notebook users intentionally cluster related operations into adjacent groups of cells to boost productivity and organization [wang2022stickyland], future research should aim to automatically generate context-aware documentation for these clusters. Utilizing analysis techniques such as dependency graph extraction and runtime tracing, researchers can identify groups of cells and automatically produce Markdown summaries.
@@
Research Direction 5: Designing Automated Bug Detection and Remediation Tools Tailored to Notebooks
@@
Since bug analysis is crucial to maintaining the quality of Jupyter notebooks (see Section 5.5), future research could focus on developing automated tools specifically designed for notebook-related bugs. Although some existing approaches address issues such as name-value inconsistencies detection [patra2022nalin] and data leakage detection [yang2022data], further research is needed to tackle more complex bugs, such as cell defects and kernel failure. Future studies should focus on addressing notebook code cell defects, which arise from missing dependencies, incorrect execution order, or undefined variables, as well as kernel crashes, often caused by excessive resource consumption, infinite loops, or faulty package interactions [de2024bug]. Advanced techniques, including static and dynamic code analysis, execution flow tracking, and visualization-based debugging, can be leveraged to efficiently detect, categorize, and resolve these bugs.
@@
Research Direction 6: Conducting Empirical Studies on the Impact of Best Practices for Coding
@@
As outlined in Section 5.7, little is known about the actual influence of following best practices for coding in Jupyter notebooks. Future studies should conduct empirical studies to quantify the effects of following best practices for coding based on categorized effective and ineffective practices [pimentel2021understanding]. Systematic studies comparing compliant and non-compliant notebooks could provide concrete empirical evidence highlighting the advantages of rigorous adherence to recommended practices. Such evidence would strongly motivate wider adoption of disciplined coding standards among notebook users.
@@
Research Direction 7: Detecting and Resolving Duplicate Execution Numbers in Notebook Cells
@@
Research Direction 8: Enhancing AI-Powered Code Generation and Workflow Assistance in Jupyter Notebook
In Section 5.9, we discuss existing solutions for generating code snippets based on natural language instructions for specific libraries like Pandas [yin2023natural]. These solutions possess limited scope with a single library and fail to support broader, multi-library, multi-step data science workflow automation. To address these limitations, future research should leverage advanced AI techniques (e.g., large language models), to improve conversational AI-powered assistants for automated code generation by multi-step workflow automation. For example, interactive low-code panels that recommend follow-up analytical questions and visualize workflow structures [chen2023whatsnext] could be refined and augmented through continuous conversational AI interactions. By dynamically adapting their recommendations based on real-time user inputs and evolving contexts, these AI-powered assistants could offer more responsive and intuitive exploratory data analysis experiences for notebook users.
@@ 7 Threats to Validity
External validity: While most studies analyzed in this review focus on Python notebooks, which may limit generalizability, this focus is justified by the fact that 93% of notebooks available on GitHub are written in Python kallen2021jupyter. This high percentage highlights the relevance of Python-based research within the broader context of Jupyter Notebook usage. Another concern regarding external validity is that our findings primarily apply to software engineering research on Jupyter notebooks. Future research should expand the scope by investigating other types of research on notebooks to improve the generalizability of the results.
Construct validity: One potential threat to the construct validity of our review is related to the study selection process. We conducted a comprehensive search of all academic articles listed in the IEEE Xplore, ACM, Springer, Elsevier, and DBLP bibliographies, intentionally excluding non-academic sources such as blog posts. Those five digital libraries are widely recognized and frequently used by software engineering researchers to identify relevant articles in systematic literature reviews.
@@ 8 Conclusion
In this systematic literature review, we provide a comprehensive overview of the research landscape of software engineering in Jupyter notebooks. Our review uncovers critical insights and trends that have shaped the field’s current state. We utilized a thorough methodology involving identifying, screening, and analyzing 199 relevant research articles (supplementary data are available here [^19] https://zenodo.org/records/17377280). The most important findings of our study are:
@@
    Most studies were published in conference venues, indicating the field is rapidly evolving. Interestingly, the most frequent venues were HCI-related instead of core SE venues. The Conference on Human Factors in Computing Systems (CHI) published 20 studies, whereas the International Conference on Software Engineering (ICSE) published 7.
@@
    Notebook-specific solutions for software engineering issues, such as testing, refactoring, and documentation, are relatively underexplored. For example, resolving duplicate execution numbers, refactoring across notebook clones, and generating group documentation for coherent code cells are future directions derived from our study.
@@
    The majority of replication packages are hosted on GitHub, raising concerns about long-term availability and adherence to open science principles, as GitHub repositories can be volatile. In contrast, only 17 studies used permanent repositories such as Zenodo or Figshare, aligning better with open science best practices.